OpenStudy (anonymous):

[f(x)][g(x)] = (x+3)(x-3) = ?

OpenStudy (anonymous):

(x+a)(x-a)=x^2-a^2

OpenStudy (anonymous):

(a+b)(a-b)=a^2+b^2

OpenStudy (mimi_x3):

x^2 - 3x + 3x - 9 x^2 - 9

OpenStudy (anonymous):

Hashir, you made a typo?